Collect like terms:
18t+t+5/3t=??
Simplify your answer . Type and improper fraction

Answer by stanbon(75887)

Collect like terms:
18t+t+5/3t=??
-----
(18 + 1 + (5/3))t
-----
= (19 + 5/3)t
---
= (19(3/3) + (5/3))t
---
= [(57/3) + (5/3)]t
---
= (62/3)t
----
= (20 2/3)t
===============
Cheers,
Stan H.